Given U=f(x,y)=Xa Yb and the budget constraint as M=PxXpyY

1
Expert's answer
2022-03-24T16:08:57-0400

Solution

Set up constrained utility maximization problem:

"Max. U=xy+x+y"

s.t "P_{x}X+P_{y}y=M"


Set up a Lagrangian function:

"L\n=\nx\ny\n+\nx\n+\ny\n\u2212\n\u03bb\n(\nP_{\nx}\nX\n+\nP_{\ny}\nY\n\u2212\nM\n)"


Take the first derivative of L with respect to x and y:

"\frac{∂
L
}{∂
x}
:
y
+
1
−
λ
P_{
x}
=
0.....
(
i
)
."


"\frac{∂
L}{
∂
y}
:
x
+
1
−
λ
P_{
y}
=
0........
(
i
i
)
."


"\frac{∂
L}{
∂
λ}
:
P_{
x}
X
+
P_{
y}
Y
−
M
=
0..........
(
i
i
i
)
."


Solve equation (i) and (ii), and solve for x and y:


"\frac{y+
1}{
P_{
x}}
=
λ
,
\frac{x
+
1}{
P_{
y}}
=
λ"


"\frac{y
+
1}{
P_{
x}}
=
\frac{x
+
1}{
P_{
y}}"


"P_{\ny}\n(\ny\n+\n1\n)\n=\nP_{\nx}\n(\nx\n+\n1\n)"


"P_{\ny}\nY\n+\nP_{\ny}\n=\nP_{\nx}\nX\n+\nP_{\nx}"


"P_{\ny}\nY\n=\nP_{\nx}\nX\n+\nP_{\nx}\n\u2212\nP_{\ny}"


"Y
=
\frac{P_{
x}
X
+
P_{
x}
−
P_{
y}}{
P_{
y}}"


"Y
=
\frac{P_{
x}
(
X
+
1
)
−
P_{
y}}{
P_{
y}}"


"P_{\ny}\nY\n\u2212\nP_{\nx}\n+\nP_{\ny}\n=\nP_{\nx}\nX"


"X
=
\frac{P_{
y}
Y
−
P_{
x}
+
P_{
y}}{
P_{
x}}"


Substitute both x and y in the budget constraint function:

"P_{
x}
X
+
P_{
y}
[\frac{
P_{
x}
(
X
+
1
)
−
P_{
y}}{
P_{
y}}
]
=
M"


"P_{\nx}\nX\n+\nP_{\nx}\n(\nX\n+\n1\n)\n\u2212\nP_{\ny}\n=\nM"


"P\nx\nX\n+\nP\nx\nX\n+\nP\nx\n\u2212\nP\ny\n=\nM"


"2\nP\nX\n+\nP\nx\n\u2212\nP\ny\n=\nM"


"2\nP\nx\nX\n=\nM\n\u2212\nP\nx\n+\nP\ny"


"X^
{M}
=\frac{
M
−
P
x
+
P
y}{
2
P
x}"


The above "X^{M}"  is the Marshallian demand for good x.


"P
x
[\frac{
P
y
Y
−
P
x
+
P
y}{
P
x
}]
+
P_{
y}
y
=
M"


"P_{\ny}\nY\n\u2212\nP_{\nx}\n+\nP_{\ny}\n+\nP_{\ny}\ny\n=\nM"


"2\nP_{\ny}\ny\n=\nM\n+\nP_{\nx}\n\u2212\nP_{\ny}"


"Y^{\nM}\n=\nM\n+\nP_{\nx}\n\u2212\nP_{\ny}\n2\nP_{\ny}"


The "Y^{M}"  is the Marshallian demand function for good Y.